OpenStudy (anonymous):

The second angle of a triangle kite is four times as large as the first. The third angle is 5 more than the sum of the other two angles. Find the measure of the second angle.

OpenStudy (anonymous):

You should know that the angle sum of a triangle is 180. Lets call the angles: a, b and c. a+b+c=180 4a=b (that is the first sentence) c-5=a+b (second sentence) Solve this system of equations to get your result.